### v1

@@ 5 Conclusion and Future Work
Acknowledgements: This work is in part supported by a gift from Adobe Research and a Richard B. Salomon award for Chen Sun. We thank helpful discussions with Professors Ellie Pavlick and Stephen Bach. We thank Yue Yang, An Yang, Yu Wang for providing the open-sourced code for the baselines.
Appendix 0.A Hyperparameters
@@
Appendix 0.B Additional Implementation Details
0.B.1 Dependency Parsing Rules
@@
0.B.2 Mutual Information Implementation
@@
Appendix 0.C Dataset Details
@@
Dataset #Class #Train #Valid #Test ImageNet 1000 128,1167 50,000 - Food-101 101 60,600 15,150 25,250 CIFAR-100 100 40,000 10,000 10,000 CIFAR-10 10 40,000 10,000 10,000 CUB-200 200 4,794 1,200 5,794 Flowers-102 102 4,093 1,633 2,463
@@
Appendix 0.D Human Evaluation Details
@@
We conduct Students’ T-test to evaluate the statistical significance of the human evaluation results. We set the threshold of p-value to be 0.05 following previous works. When p-value is lower than 0.05, the null hypothesis is rejected and out method performs significantly better than the baseline method. From the results we can observe that both our concept learning and concept discovery method significantly outperform the baseline methods regarding the precision and thoroughness metrics.
@@
Appendix 0.E Ablation Study on Effectiveness of CDL Framework
@@
The results in Table A3 and A4 indicate that (1) it is the concept discovery that mainly contribute to the improvement of classification performance and the concept learning would not affect the classification accuracy; (2) both concept discovery and learning parts provide significant improvement for the interpretability of the concepts according to the intervention accuracy resutls.
@@
Appendix 0.F Few-shot Performance about Generalization

### v2

@@ 5 Conclusion and Future Work
6 Acknowledgement
This work is in part supported by a gift from Adobe Research, a seed grant from NASA, and a Richard B. Salomon award for Chen Sun. We thank helpful discussions with Professors Ellie Pavlick and Stephen Bach. We thank Yue Yang, An Yang, Yu Wang for providing the open-sourced code for the baselines.
Appendix A Hyperparameters
@@
Appendix B Additional Implementation Details
B.1 Dependency Parsing Rules
@@
B.2 Mutual Information Implementation
@@
Appendix C Dataset Details
@@
Dataset #Class #Train #Valid #Test ImageNet 1000 128,1167 50,000 - Food-101 101 60,600 15,150 25,250 CIFAR-100 100 40,000 10,000 10,000 CIFAR-10 10 40,000 10,000 10,000 CUB-200 200 4,794 1,200 5,794 Flowers-102 102 4,093 1,633 2,463 Stanford Cars 196 6,515 1,628 8041 Aircrafts 102 3,400 3,400 3,400 Oxford Pets 37 1,846 1,834 3669
@@
Appendix D Human Evaluation Details
@@
Appendix E Statistical Significance
We conduct Students’ T-test to evaluate the statistical significance of our main experiments including classification accuracy and human evaluation of concept quality. We set the threshold of p-value to be 0.05 following previous works. When p-value is lower than 0.05, the null hypothesis is rejected and out method performs significantly better than the baseline method. From the results in Table A2 we can observe that both our proposed method significantly outperform the baseline methods regarding the classification performance and concept quality.
ImageNet Food CIFAR-100 CIFAR-10 CUB Flowers Stanford Cars Aircrafts Oxford Pets CDL v.s. LaBo 9e-02 (✗) 2e-74 (✓) 1e-09 (✓) 3e-14 (✓) 1e-06 (✓) 3e-03 (✓) 1e-02 (✓) 2e-01 (✗) 1e-02 (✓) CDL v.s. LM4CV 4e-01 (✗) 1e-68 (✓) 9e-05 (✓) 2e-03 (✓) 1e-14 (✓) 2e-01 (✗) 3e-02 (✓) 4e-01 (✗) 3e-02 (✓)
Table A3: The statistical significance of the classification accuracy. Smaller p-value indicates higher significance. The performance gain is significant when p-value is lower than 5e-02.
Appendix F Ablation Study on Effectiveness of CDL Framework
@@
The results in Table A4, A5 and A6 indicate that (1) it is the concept discovery that mainly contribute to the improvement of classification performance and the concept learning would not affect the classification accuracy; (2) both concept discovery and learning parts provide significant improvement for the interpretability of the concepts according to the intervention accuracy resutls.
@@
Table A6: Human evaluation of the discovered concepts before and after concept learning on ImageNet.
Appendix G Few-shot Performance about Generalization
@@
Appendix H Broader Application of Visual Concepts
Besides image classification tasks, the discovered and learned visual concepts can also benefit tasks that require object localization by localizing visual concepts in images. We integrate the visual concepts into the existing object detection framework OWL proposed by Minderer et al. (2022) and conduct evaluations on the MS-COCO dataset (Lin et al., 2014). We first perform concept discovery and learning on the training data, and then replace the category names (e.g. “dog”) with selected visual concepts (e.g. “four-legged mammal” and "floppy ears") to train and evaluate the object detection model. The results are shown in Table A7. The visual concepts can enhance the object detection performance of the existing framework.
Method AP AP50 OWT 43.5 64.7 OWT + CDL 46.2 68.1
Table A7: Object detection performance on MS-COCO dataset. The baseline results are quoted from Minderer et al. (2022). Visual concept discovery and learning can enhance the object detection performance of the existing framework.